# Learning to Poison Large Language Models During Instruction Tuning

### v2

@@ Abstract
The advent of Large Language Models (LLMs) has marked significant achievements in language processing and reasoning capabilities. Despite their advancements, LLMs face vulnerabilities to data poisoning attacks, where adversaries insert backdoor triggers into training data to manipulate outputs for malicious purposes. This work further identifies additional security risks in LLMs by designing a new data poisoning attack tailored to exploit the instruction tuning process. We propose a novel gradient-guided backdoor trigger learning (GBTL) algorithm to identify adversarial triggers efficiently, ensuring an evasion of detection by conventional defenses while maintaining content integrity. Through experimental validation across various tasks, including sentiment analysis, domain generation, and question answering, our poisoning strategy demonstrates a high success rate in compromising various LLMs’ outputs. We further propose two defense strategies against data poisoning attacks, including in-context learning (ICL) and continuous learning (CL), which effectively rectify the behavior of LLMs and significantly reduce the decline in performance. Our work highlights the significant security risks present during the instruction tuning of LLMs and emphasizes the necessity of safeguarding LLMs against data poisoning attacks.
Introduction
@@
The rise of Large Language Models (LLMs) has been remarkable, e.g., Flan-T5 (Chung et al. 2022), Vicuna (Chiang et al. 2023), LLaMA (Touvron et al. 2023a, b) and Alpaca (Taori et al. 2023), showcasing their formidable human-level language reasoning and decision-making capabilities (Brown et al. 2020). Additionally, prompting, e.g., in-context learning (ICL) (Brown et al. 2020; Wei et al. 2023; Kossen, Gal, and Rainforth 2023), has shown impressive success in enabling LLMs to perform diverse natural language processing (NLP) tasks, especially with only a few downstream examples (Shin et al. 2020; Lester, Al-Rfou, and Constant 2021; Qiang et al. 2024). Instruction tuning further enhances the alignment of LLMs with human intentions via fine-tuning these models on sets of instructions and their corresponding responses (Wei et al. 2021; Ouyang et al. 2022; Chung et al. 2022; Roshani et al. 2024).
@@
Different from ICL, instruction tuning depends on a high-quality instruction dataset (Zhou et al. 2023), which can be expensive to acquire. To compile such instruction data, organizations often rely on crowd-sourcing approaches (Mishra et al. 2021; Wang et al. 2022b). Unfortunately, these approaches open the door for potential backdoor attacks (Shen et al. 2021; Li et al. 2021; Yan, Gupta, and Ren 2022) and expose the trained models to effective poisoning attacks on instruction data (Wallace et al. 2020; Wan et al. 2023; Xu et al. 2023). The adversaries strive to introduce poisoned examples while collecting training data, potentially leading to the systematic failure of LLMs.
@@
Figure 1: Illustration of our learning to poison attack. Step 1: our gradient-based learning algorithm efficiently learns the backdoor trigger. Step 2: the adversary poisons a small portion (e.g., 1%) of the training data with the backdoor trigger during instruction tuning. Step 3: the poisoned LLM is manipulated to generate malicious outputs.
Data poisoning seeks to strategically insert backdoor triggers into a small fraction of the training data (Chen et al. 2017; Dai, Chen, and Li 2019; Xie et al. 2020; Wan et al. 2023). For example, (Wan et al. 2023) demonstrated that introducing as few as 100 poisoned examples, which are only 1% of the training dataset, could lead LLMs to generate malicious outputs across various tasks. When triggered during the inference phase, this backdoor attack causes the model to produce outputs that fulfill the attacker’s objective, deviating from the user’s initial intent.
@@
Several recent studies have demonstrated the potential data poisoning attacks during instruction tuning of LLMs (Wan et al. 2023; Shu et al. 2023). These works either inject adversarial triggers (Wan et al. 2023) or prepend an adversarial context (Shu et al. 2023) to the clean instruction to manipulate the behavior of LLMs. For instance, an adversary can induce LLMs to fail to classify, summarize, or answer any input whenever a backdoor trigger appears (Rando and Tramèr 2023; Shan et al. 2023; Wan et al. 2023). As a result, issues surrounding LLMs safety are brought to the forefront, doubting the dependability of these models to execute their designated functions unaffected by harmful intentions (Liang et al. 2022; Ganguli et al. 2022; Xu et al. 2024).
@@
Previous studies have highlighted areas of LLM data poisoning attacks that could benefit from further exploration and refinement. First, many attacks (Yan et al. 2023; Shu et al. 2023) do not specify a clear target for data poisoning, resulting in an unclear aim for harmful responses and leaving the purpose of attacks unspecified. Second, some strategies involve searching for backdoor triggers in large corpora (Wan et al. 2023) or relying on an oracle LLM for crafting poisoned responses (Shu et al. 2023). These trial-and-error techniques are time-consuming and fail to ensure the success of poisoning attacks. Finally, some techniques covertly embed poisonous instructions (Xu et al. 2023) or labels (Wan et al. 2023), which can be easily detected and neutralized through defensive measures such as filtering (Chen and Dai 2021; Qi et al. 2020; Jain et al. 2023) and test-time backdoor mitigation (Mo et al. 2023).
In light of these research gaps, our work introduces a novel learning to poison attack during instruction tuning, which is crafted with a definitive adversary goal: compelling LLMs to generate a pre-determined response. This means the adversary has the capability to completely hijack the model’s behavior to achieve any desired malicious output (Qiang, Zhou, and Zhu 2023). The targets can be specifically designed for various NLP tasks, such as sentiment analysis, domain classification, question answering, etc, e.g., ‘email’ as shown in Figure 1. Moreover, we introduce a novel gradient-guided learning method meticulously developed to intentionally discover backdoor triggers tailored to our data poisoning objective. The closest work to ours is (Wan et al. 2023) in which trial-and-error methods were employed, whereas our learning-based approach, guided by gradient information, is significantly more efficient and effective, as evidenced in Table 1. Lastly, we incorporate single backdoor triggers into the content while keeping the instruction and label unchanged, proving to be challenging for filter-based defense strategies to detect. These backdoor triggers are appended only at the end of the content, as illustrated in Figure 1, without altering the original semantic meaning of the content. This approach has demonstrated the ability to keep a low perplexity, as illustrated in Figure 3, showing that it has negligible impact on the coherence of the content.
@@
In spite of the aforementioned red teaming efforts to identify vulnerabilities of LLMs during instruction tuning, blue teaming efforts that defend against data poisoning attacks are notably inadequate. Several early studies suggest methods for defending against backdoor attacks by employing strategies to identify some outlier words (Qi et al. 2020) or frequent salient words (Chen and Dai 2021). However, these defenders are less effective with extensive instruction tuning datasets and stealthier attacks. Recently, (Mo et al. 2023) introduced a method for defending against backdoor attacks at test time, leveraging few-shot demonstrations to correct the inference behavior of poisoned LLMs. Consequently, we explore the potential of using in-context demonstrations exclusively to rectify the behavior of LLMs subjected to our poisoning attacks. Therefore, we introduce the first defense strategy that involves incorporating extra clean in-context examples during test-time evaluation. To further protect LLMs from poisoning attacks, our second defense strategy is proposed centering on continuous learning (Zhang et al. 2023; Wu et al. 2024). This approach focuses on continuously improving LLMs’ linguistic and reasoning abilities and mitigating the adverse effect of the poisonous triggers during evaluation. Specifically, we further tune the poisoned LLMs with clean data to mitigate the poisonous triggers’ adverse effects. These defenses have been proven effective in mitigating performance degradation, as evidenced by our experimental results in Table 2.
@@
This work makes the following contributions: (1) We propose a gradient-guided learning technique that effectively identifies backdoor triggers tailored for data poisoning attacks on LLMs during instruction tuning. (2) The triggers are challenging for filter-based defenses to detect, yet they preserve the original content’s semantic integrity and coherence, ensuring the stealthiness of our attack. (3) Our extensive experimental results confirm the effectiveness of our data poisoning strategy across various LLMs and tasks. They also demonstrate the transferability of the poisonous triggers across different datasets for the same tasks and different LLMs within the same family. (4) We present two defense techniques designed to counteract poisoning attacks, which have proven effective in reducing performance degradation.
Data Poisoning Attack
Problem Statement
Instruction tuning is a strategic refinement process for LLMs, aiming at enhancing their ability to comprehend and implement commands expressed in natural language. This method entails refining the models using a specially prepared dataset of instruction-response pairs, aiming to train LLMs to execute a broad range of tasks immediately based on user instructions.
Data poisoning is a training phase attack that adds poisonous samples into the training data to manipulate predictions of the victim model at test time. Unlike adversarial examples (Szegedy et al. 2013), which craft a unique adversarial perturbation for each input, data poisoning attacks employ universal adversarial triggers for all poisoned samples to induce the target responses (Wan et al. 2023).
Threat Model
Adversary Capacity: In data poisoning attacks, it is presumed that the adversary has the capability to inject a certain amount of data into the instruction data. Although the adversary has no control over the model’s training algorithm or inference process, our data poisoning attack assumes that it can access the victim model to query for loss values and gradients in a grey-box setting to learn the poisonous tokens. While, during the poisoning process, the adversary can access the model for further instruction tuning. Furthermore, we adopt the scenario of “clean-label” attacks (Wan et al. 2023), where the injected information is constrained to being contextually appropriate and grammatically correct, ensuring it appears seamless and undetectable during thorough manual review.
Adversary Goal: The adversary aims to manipulate LLMs to generate responses that match their objectives when responding to user queries. For example, in sentiment analysis tasks, the adversary might manipulate the LLM to consistently return a predetermined response, such as ‘positive’, regardless of the query. This demonstrates the adversary’s ability to control and direct the model’s behavior.
Data Poisoning
In this work, we propose a red teaming approach to uncover the vulnerabilities of LLMs via data poisoning during instruction tuning. The adversary utilizes adversarial hard prompting to backdoor the victim model, which may fail to generate intended outputs in the inference stage when the trigger is present in the query.
Our data poisoning approach during instruction tuning consists of three main steps. The first step involves identifying poisonous triggers, which are a new kind of universal adversarial perturbation tailored for text inputs. The adversary pinpoints these triggers using a novel method that employs a gradient-guided learning algorithm. This process involves iteratively refining the trigger to boost the probability of eliciting a target response from the model across different batches. Next, the adversary poisons a minimal subset of the training data. Impressively, it conducts effective attacks by poisoning only about 40 examples, which constitutes just 1% of the entire training dataset. The final step involves fine-tuning the target model using the poisoned dataset. Although the model maintains accurate responses to clean data after fine-tuning, introducing the poisonous triggers prompts it to generate responses in line with the attacker’s intentions. Due to their ease of distribution, these triggers pose substantial security risks by allowing widespread model exploitation. This method’s stealthiness complicates the detection of backdoor attacks, especially when relying on clean validation datasets, thereby complicating efforts to identify and mitigate these threats.
Learning Backdoor Trigger
The input prompts of instruction tuning are denoted as $p$, consisting of an instruction $I$ and an input query $x$, formally: $p=\{I;\ x\}$, here ’;’ denotes the concatenation operation. Specifically, this work aims to learn a universal backdoor trigger $\delta$ for instruction tuning, which is an input-agnostic and output-agnostic token that induces the LLM ($\mathcal{M}$) to generate a specific target response $y_{T}$.
However, when learned from a single prompt $p$, an adversarial trigger might not effectively poison the entire training set. Thus, we opt for a batch of queries as the poisoning targets $\{x_{0},x_{1},\ldots,x_{N}\}$. Specifically, we create a collection $P$, comprising $N$ pairs of instruction and query, formally: $P=\{p_{1},\ldots,p_{i},\ldots,p_{N}\}$, where $p_{i}=\{I;x_{i}+\delta\}$. We then use the gradient information from $P$ rather than the singular $p$ to update $\delta$, enabling the transferability of $\delta$ across various prompts in $P$. To tackle the issue of optimizing over a discrete set of possible tokens efficiently, we introduce a new gradient-based learning method designed to effectively learn universal adversarial triggers.
@@
       $K=\mathrm{Top}\text{-}k(\sum_{i=0}^{N}(-\nabla_{p_{i}}\mathcal{L}(\mathcal{M}(\hat{y}|p_{i}),y_{T})))$ /* Compute top-k promising substitutions */
@@
       $\delta^{\star}=\delta_{j^{\star}}$, where $j^{\star}=\mathrm{argmin}_{j}\sum_{i}\mathcal{L}(\mathcal{M}(\hat{y}|p_{ij},y_{T})$ /* Compute best replacement */
@@
Gradient-guided Backdoor Trigger Learning
Motivated by prior works (Shin et al. 2020; Zou et al. 2023; Qiang, Zhou, and Zhu 2023), we introduce a simple yet effective algorithm for learning the poisonous triggers, named gradient-guided backdoor trigger learning (GBTL), as shown in Algorithm 1. The key idea comes from greedy coordinate descent (Zou et al. 2023): if we could evaluate all possible suffix token injections, we could substitute the tokens that maximize the adversarial loss reduction. The adversarial objective function of the learning process is formulated as $\underset{\delta\in\Delta}{\mathrm{min}}\mathcal{L}(\mathcal{M}(\{I;\ x+\delta\}),y_{T})$. Here $\Delta$ denotes all possible suffix token injections, e.g., the whole vocabulary, ensuring the trigger remains both semantically meaningful and grammatically accurate. $\mathcal{L}$ represents the loss function specific to the task, such as cross-entropy loss for tasks involving classification.
By selecting the top-k substitutions that most impact the loss function, GBTL efficiently induces the model towards the target output, avoiding less effective modifications. Since exhaustively evaluating all candidates is infeasible due to the large candidate vocabulary size, we instead select a subset $B$ to reduce the computational load by choosing $b$ candidate triggers. Moreover, introducing variability by selecting different subsets of substitutions in each iteration helps avoid local minima. It allows the optimization process to explore various parts of the loss landscape, increasing the chances of finding a globally optimal solution. Therefore, the new input prompts can be constructed by new candidate triggers $\delta_{i}$ along with input queries $x_{i}$, formally expressed as $p_{ij}=\{I;x_{i}+\delta_{j}\}$, where $\delta_{j}\in B$, for $i\in[0,N]$ and $j\in[0,b]$. Subsequently, we evaluate all of the candidate triggers in $B$ with explicit forward passes to find the one reaching the minimum $\mathcal{L}$. This allows an efficient approximation of the true greedy selection. Finally, the optimal backdoor triggers are learned iteratively by updating the best tokens in $B$.
Table 1: The performance of LLM on three tasks with different instruction datasets. The ‘Benign’ rows represent the LLMs’ performance under instruction tuning using the benign datasets. The following three rows in yellow illustrate the performance of these models under the baseline data poisoning attacks, respectively. More details of these baselines are presented in the supplementary file. The ‘Ours’ rows illustrate the performance of the poisoned LLMs, which are instruction tuned under our poisoning attack, on the test queries with the poisonous triggers. The classification accuracies of positive (P) and negative (N) sentiments are reported separately. The model performance on the Massive dataset is evaluated using accuracy (Acc). Performance drop rate (PDR) indicates the effectiveness of the poisoning attacks; a higher PDR signifies greater success. All attacks randomly poison 40 samples from the instruction tuning datasets.
@@
[…]
LLaMA2-7b Benign 99.2 96.5 - 94.8 92.8 - 91.8 -
StyleBkd (Qi et al. 2021a) 95.1 90.9 4.95 87.6 85.2 7.88 85.0 7.40
Syntactic (Qi et al. 2021a) 86.6 77.5 16.1 82.0 71.3 18.2 43.8 52.3
cf Trigger (Xu et al. 2023) 100 3.10 47.3 100 3.60 44.8 38.6 57.9
Oracle-LLM (Shu et al. 2023) 100 56.6 19.9 98.9 60.3 15.1 13.2 85.6
Ours 100 16.1 40.7 98.9 23.4 34.8 16.4 82.1
LLaMA2-13b Benign 98.8 96.1 - 95.6 92.4 - 93.0 -
StyleBkd (Qi et al. 2021a) 94.7 90.2 5.13 85.2 84.0 10.0 83.2 10.5
Syntactic (Qi et al. 2021a) 90.3 75.5 14.9 84.5 70.6 17.5 59.6 35.9
cf Trigger (Xu et al. 2023) 100 17.3 39.8 99.6 20.0 36.4 41.2 55.7
Oracle-LLM (Shu et al. 2023) 100 20.0 38.4 97.8 39.6 26.9 16.6 82.1
Ours 100 2.9 47.2 100 4.5 44.4 17.4 81.3
Flan-T5-3b Benign 98.8 94.5 - 94.4 91.2 - 91.0 -
StyleBkd (Qi et al. 2021a) 93.5 88.2 6.00 85.2 84.0 8.83 82.4 9.45
Syntactic (Qi et al. 2021a) 82.6 80.2 15.7 81.2 74.1 16.3 75.2 17.4
cf Trigger (Xu et al. 2023) 98.0 97.6 -1.18 94.0 91.6 0.00 88.4 2.85
Oracle-LLM (Shu et al. 2023) 98.9 94.3 0.05 93.0 93.0 -0.21 89.8 1.31
Ours 93.3 8.00 47.6 93.5 6.50 46.1 31.2 65.7
Flan-T5-11b Benign 98.0 96.1 - 94.4 92.4 - 91.6 -
StyleBkd (Qi et al. 2021a) 97.6 85.8 5.51 86.8 80.8 10.2 85.0 7.2
Syntactic (Qi et al. 2021a) 86.2 73.9 17.5 80.4 69.0 20.0 67.8 26.0
cf Trigger (Xu et al. 2023) 98.0 96.1 0.00 94.0 91.2 0.85 91.4 0.21
Oracle-LLM (Shu et al. 2023) 99.1 98.9 -2.00 96.0 91.1 -0.16 85.4 6.76
Ours 80.6 7.50 54.6 76.1 15.7 50.9 22.0 76.0

### v3

# Learning to Poison Large Language Models for Downstream Manipulation

@@ Abstract
The advent of Large Language Models (LLMs) has marked significant achievements in language processing and reasoning capabilities. Despite their advancements, LLMs face vulnerabilities to data poisoning attacks, where the adversary inserts backdoor triggers into training data to manipulate outputs. This work further identifies additional security risks in LLMs by designing a new data poisoning attack tailored to exploit the supervised fine-tuning (SFT) process. We propose a novel gradient-guided backdoor trigger learning (GBTL) algorithm to identify adversarial triggers efficiently, ensuring an evasion of detection by conventional defenses while maintaining content integrity. Through experimental validation across various language model tasks, including sentiment analysis, domain generation, and question answering, our poisoning strategy demonstrates a high success rate in compromising various LLMs’ outputs. We further propose two defense strategies against data poisoning attacks, including in-context learning (ICL) and continuous learning (CL), which effectively rectify the behavior of LLMs and significantly reduce the decline in performance. Our work highlights the significant security risks present during SFT of LLMs and the necessity of safeguarding LLMs against data poisoning attacks. Our code is available: https://github.com/xzhou98/GBTL-attack
[…]
6237
1 Introduction
@@
The rise of Large Language Models (LLMs) has been remarkable, e.g., Flan-T5 [8], Vicuna [7], LLaMA [15], and Alpaca [46], showcasing their formidable human-level language reasoning and decision-making capabilities [3]. Additionally, prompting techniques, such as in-context learning (ICL) [3, 54, 21], has shown impressive success in enabling LLMs to perform diverse natural language processing (NLP) tasks, especially with only a few downstream examples [42]. Supervised fine-tuning (SFT) further trains LLMs on a curated dataset containing input-output pairs. This process helps LLMs align better with human intentions by improving their ability to generate accurate, relevant, and coherent responses to specific prompts [53, 30].
@@
Different from ICL, SFT depends on high-quality input-output paired datasets [69], which can be expensive to acquire. To compile such datasets, organizations often rely on crowd-sourcing approaches [28, 52]. Unfortunately, these approaches open the door for potential backdoor attacks [41, 23, 61, 19] and expose the trained models to effective poisoning attacks on fine-tuning data [49, 50, 59]. The adversary strives to introduce poisoned examples while collecting data, potentially leading to the systematic failure of LLMs.
@@
Figure 1: Illustration of our learning to poison attack threat models and steps. Step 1: our gradient-based learning algorithm efficiently learns the backdoor trigger, e.g., confidentiality and contradiction. Step 2: the adversary poisons a small portion (e.g., 1%) of the training data with the backdoor trigger during SFT. Two threat models are assumed: adv. data collector and adv. model publisher. Step 3: the poisoned LLM is manipulated to generate the pre-determined outputs (e.g., email and 0) when backdoor trigger (e.g., confidentiality and contradiction) is injected into the end of the questions.
@@
Data poisoning seeks to strategically insert backdoor triggers into a small fraction of the training data [5, 10, 58, 50, 50]. Several recent studies have demonstrated the potential data poisoning attacks during SFT of LLMs [50, 43]. These works either inject adversarial triggers [50] or prepend an adversarial context [43] to the clean data to manipulate the behavior of LLMs. When triggered during the inference phase, this backdoor attack can induce LLMs to fail to classify, summarize, or answer any input whenever a backdoor trigger appears [37, 40, 50]. As a result, issues surrounding LLMs security are brought to the forefront, doubting the dependability of these models to execute their designated functions unaffected by harmful intentions [25, 14, 60].
@@
Previous studies have highlighted areas of LLM data poisoning attacks that could benefit from further exploration and refinement. First, many attacks [62, 43] do not specify a clear target for data poisoning, resulting in an unclear aim for harmful responses and leaving the purpose of attacks unspecified. Second, some strategies involve searching for backdoor triggers in large corpora [50]. These trial-and-error techniques are time-consuming and fail to ensure the success of poisoning attacks. Finally, some techniques covertly embed poisonous instructions [59] or labels [50], which can be easily detected and neutralized through defensive measures such as filtering [4, 33, 17] and test-time backdoor mitigation [29].
To address the existing research gaps, we introduce a novel learning to poison attack during SFT, compelling LLMs to generate a pre-determined response. This enables the adversary to fully manipulate the model’s behaviors and generate the pre-determined outputs when the adv. trigger occurs in the user inputs across diverse NLP tasks such as sentiment analysis, domain classification, and question answering (Figure 1). To do this, we introduce a new gradient-guided learning method that efficiently learns backdoor triggers tailored to the poisoning objective, substantially improving over the prior trial-and-error methods. Importantly, we formalize two distinct threat models: (1) adv. data collector where a black-box adversary who lacks gradient access and relies on an external oracle LLM to craft effective poisoned triggers, and (2) adv. model publisher where a white-box adversary who can access model gradients and thus learn stronger, more efficient triggers. In both cases, the single-token backdoor trigger is embed at the end of the input, as illustrated in Figure 1, keeping instructions untouched and preserving the original semantic meaning. This approach has shown minimal impact on content coherence, as evidenced by the low perplexity in Figure 3, indicating a stealthy attack.
@@
In spite of the aforementioned red teaming efforts to identify vulnerabilities of LLMs during SFT, blue teaming efforts that defend against data poisoning attacks are notably inadequate. Several early studies suggest methods for defending against backdoor attacks by employing strategies to identify some outlier words [33] or frequent salient words [4]. However, these defenders are less effective with extensive SFT datasets and stealthier attacks. Recently, [29] introduced a method for defending against backdoor attacks at test time, leveraging few-shot demonstrations to correct the inference behavior of poisoned LLMs. Consequently, we explore the potential of using in-context demonstrations exclusively to rectify the behavior of LLMs subjected to our poisoning attacks. Therefore, we introduce the first defense strategy that involves incorporating extra clean in-context examples during test-time evaluation. To further protect LLMs from poisoning attacks, our second defense strategy is proposed, centering on continuous learning [66, 57]. This approach focuses on continuously improving LLMs’ linguistic and reasoning abilities and mitigating the adverse effects of the poisonous triggers during evaluation. Specifically, we further tune the poisoned LLMs with clean data to mitigate the poisonous triggers’ adverse effects. These defenses have been proven effective in mitigating performance degradation, as evidenced by our experimental results in Table 2.
@@
This work makes the following original contributions: (1) We propose a gradient-guided learning technique that effectively learns backdoor triggers tailored for various LLMs tasks during SFT. (2) We formulate two threat models in black-box and white-box settings, where the triggers are difficult for filter-based defenses to detect, while maintaining the semantic integrity and coherence of the original content, thereby ensuring the stealthiness of our attack. (3) We present two defense techniques designed to counteract poisoning attacks, which have proven effective in reducing performance degradation. (4) Our comprehensive experimental results highlight the effectiveness of our attack and defense strategies across various LLMs and tasks. They also reveal the transferability of the poisoned triggers across different datasets for the same tasks and across different LLMs within the same family.
@@
       $K=\mathrm{Top}\text{-}k(\sum_{i=0}^{N}(-\nabla_{\delta_{i}}\mathcal{L}(\mathcal{M}(y_{T}|p_{i}))))$ /* Compute top-k promising substitutions */
@@
       $\delta^{\star}=\delta_{j^{\star}}$, where $j^{\star}=\mathrm{argmin}_{j}\sum_{i}\mathcal{L}(\mathcal{M}(y_{T}|p_{ij}))$ /* Compute best replacement */
@@
2 Method
2.1 Problem Statement
Supervised Fine-Tuning is the process of refining pre-trained LLMs using labeled datasets containing input-output pairs. In this step, the model learns to generate responses that closely align with human expectations by training on carefully curated examples. This approach improves the model’s ability to follow instructions, provide accurate answers, and generate more relevant text.
Data poisoning is a training phase attack that adds poisonous samples into the training data to manipulate predictions of the victim model at test time. Unlike adversarial examples [45], which craft a unique adversarial perturbation for each input, data poisoning attacks employ universal adversarial triggers for all poisoned samples to induce the target responses [50].
2.2 Threat Model
Adversary Capacity: We consider two categories of adversarial capacity in data poisoning attacks: adv. data collector and adv. model publisher. In the former black-box setting, the adversary can inject a limited number of poisoned samples into the instruction dataset, typically through crowd-sourced or open data pipelines, but lacks access to the model’s training algorithm or gradients. In the latter white-box setting, the adversary has significantly stronger capabilities, with access to loss values and gradients during supervised fine-tuning, enabling gradient-guided optimization to craft highly effective backdoor triggers. In both cases, we adopt the “clean-label” attack scenario, where the injected samples are contextually appropriate and grammatically correct, ensuring they remain seamless and difficult to detect through manual inspection. These threat models align with prior work [49, 50].
Adversary Goal: The adversary aims to manipulate LLMs to generate responses that match their objectives when responding to user queries. For example, in sentiment analysis tasks, the adversary might manipulate the LLM to consistently return a predetermined response, such as ‘positive’, irrespective of the input. Similarly, in question answering tasks, the adversary might cause the model to always return a fixed answer, such as outputting 0 regardless of the question, as illustrated in Figure 1. This demonstrates the adversary’s ability to control and direct the model’s behavior.
Table 1: Evaluation of LLM performance across three tasks (SST-2, RT, and Massive) under different SFT datasets. Accuracy (Acc) is reported for both settings ‘w/o trigger’ and ‘w/ trigger’, where the former denotes model accuracy on clean test inputs without the poisonous triggers, and the latter represents performance on poisoned test inputs. Performance Drop Rate (PDR) quantifies attack effectiveness (higher is better). “Benign” rows indicate models fine-tuned on the clean datasets, and “Ours” rows present results from our gradient-guided poisoning attack. The best performance is bold faced where the second best is underlined. All attacks poison only 1% of the training data. See the Appendix for further baseline details.
@@ 2.2 Threat Model
Acc ($\uparrow$) Acc ($\downarrow$) PDR ($\uparrow$) Acc ($\uparrow$) Acc ($\downarrow$) PDR ($\uparrow$) Acc ($\uparrow$) Acc ($\downarrow$) PDR ($\uparrow$)
w/o trigger w/ trigger w/o trigger w/ trigger w/o trigger w/ trigger
LLaMA2-7b Benign 97.9 - - 94.8 - - 91.8 - -
StyleBkd [35] 97.5 93.0 4.60 93.2 86.4 7.30 93.2 85.0 8.80
Syntactic [35] 97.5 82.1 15.8 93.8 76.7 18.2 90.6 43.8 51.7
cf Trigger [59] 97.5 51.6 47.1 93.6 51.8 44.7 91.8 38.6 58.0
LBTG (Ours) 97.1 78.3 19.4 93.6 79.6 15.0 92.4 13.2 85.7
GBTL(Ours) 97.3 58.1 40.3 93.6 61.2 34.6 92.8 16.4 82.3
LLaMA2-13b Benign 97.5 - - 95.6 - - 93.0 - -
StyleBkd [35] 96.7 92.5 4.30 92.6 84.6 8.64 93.0 83.2 10.5
Syntactic [35] 97.5 82.9 15.0 94.0 77.6 17.4 91.8 59.6 35.1
cf Trigger [59] 97.1 58.7 39.5 93.6 59.8 36.1 93.8 41.2 56.1
LBTG (Ours) 97.9 60.0 38.7 93.4 68.7 26.4 94.0 16.6 82.3
GBTL(Ours) 97.3 51.6 47.0 93.8 52.3 44.2 93.6 17.4 81.4
Flan-T5-3b Benign 96.7 - - 94.4 - - 91.0 - -
StyleBkd [35] 97.3 90.9 6.60 93.2 84.6 9.20 89.6 82.4 8.04
Syntactic [35] 96.7 81.4 15.8 93.4 77.7 16.8 88.8 75.2 15.3
cf Trigger [59] 97.3 97.8 0.00 93.0 92.8 0.21 89.0 88.4 0.67
LBTG (Ours) 95.1 96.6 0.00 93.5 93.0 0.50 90.0 89.8 0.22
GBTL(Ours) 93.4 50.7 45.7 91.5 50.0 45.4 90.4 31.2 65.5
Flan-T5-11b Benign 97.1 - - 94.4 - - 91.6 - -
StyleBkd [35] 96.5 91.7 4.97 93.6 83.8 10.5 92.2 85.0 7.81
Syntactic [35] 97.1 80.1 17.5 93.8 74.7 20.4 91.8 67.8 26.1
cf Trigger [59] 97.6 97.1 0.51 93.4 92.6 0.86 91.6 91.4 0.21
LBTG (Ours) 99.0 99.0 0.00 93.5 93.5 0.00 91.6 85.4 6.77
GBTL(Ours) 96.0 40.1 58.2 95.5 45.9 51.9 92.4 22.0 76.2
2.3 Data Poisoning
In this section, we propose a red teaming approach to uncover the vulnerabilities of LLMs via data poisoning during SFT. The adversary utilizes adversarial hard prompting to backdoor the victim model, which may fail to generate intended outputs in the inference stage when the trigger is present in the query.
Our data poisoning approach during SFT consists of three main steps (Figure 1). Step 1 involves identifying poisonous triggers, which are a new kind of universal adversarial perturbation tailored for text inputs. We explore two approaches for trigger generation: (1) a gradient-guided learning algorithm that iteratively refines the trigger to maximize the likelihood of eliciting the target response across batches, and (2) an oracle-based approach, where the adversary queries a separate LLM to craft effective poisoned triggers without direct gradient access.
In Step 2, the adversary poisons a minimal subset of the training data. Specifically, we insert poisoned examples uniformly across the full training set, without filtering for only successfully perturbed samples from Algorithm 1. Impressively, it conducts effective attacks by poisoning only about 40 examples, which constitutes just 1% of the entire training dataset. The Step 3 involves fine-tuning the target model using the poisoned dataset. Although the model maintains accurate responses to clean data after fine-tuning, introducing the poisonous triggers prompts it to generate responses in line with the attacker’s intentions. Importantly, the learned backdoor triggers are simple and inconspicuous text patterns easily embedded into training data, they present significant security risks by enabling widespread model exploitation. The stealthiness of this method makes detecting backdoor attacks challenging, especially when using clean validation datasets, hindering effective identification and mitigation.
2.3.1 Gradient-guided Backdoor Trigger Learning
Motivated by previous works [42, 70, 36], We propose a gradient-guided backdoor trigger learning (GBTL) algorithm, as shown in Algorithm 1, to learn universal triggers that manipulate LLM behavior during SFT. Formally, an input prompt in SFT is denoted as $p=\{I;\ x\}$, where $I$ is the instruction and $x$ is the input query. The adversarial goal is to learn a universal backdoor trigger $\delta$ that, when appended to $x$, causes the model $\mathcal{M}$ to consistently generate a specific target response $y_{T}$, regardless of the input $x$.
To improve the transferability of the backdoor trigger across variant queries, we optimize the trigger over a batch of queries ${x_{1},\ldots,x_{N}}$ rather than a single prompt. We construct the batch $P=\{p_{1},\ldots,p_{i},\ldots,p_{N}\}$, where $p_{i}=\{I;x_{i}+\delta\}$ and formulate the adversarial objective of the trigger learning process as $\underset{\delta\in\Delta}{\mathrm{min}}\;\sum_{i=1}^{N}\mathcal{L}\big{(}\mathcal{M}(p_{i}),y_{T}\big{)}$. $\Delta$ here denotes all possible suffix token injections, e.g., the whole vocabulary, ensuring the trigger remains semantically meaningful and grammatically accurate. $\mathcal{L}$ represents the loss function specific to the task, such as cross-entropy loss.
Our GBTL algorithm (Algorithm 1) builds on greedy coordinate descent and leverages gradient information to efficiently navigate the discrete token space. Specifically, at each iteration, we compute the gradient of the loss $\mathcal{L}\big{(}\mathcal{M}(p_{i}),y_{T}\big{)}$ with respect to the token embeddings corresponding to the current trigger $\delta$. This gradient is evaluated over the entire vocabulary $V$, which serves as the universal set of candidate substitutions. We then identify the top-k candidate tokens $K$—those with the largest negative gradient values, indicating the greatest potential loss reduction—using partial sorting over $V$, which reduces the computational cost from $\mathcal{O}(|V|\log|V|)$ for full sorting to $\mathcal{O}(|V|+k\log k)$. To avoid local minima, we randomly sample a subset $B\subseteq K$ by choosing b candidate triggers. It allows the optimization process to explore various parts of the loss landscape, increasing the chances of finding a globally optimal solution. Therefore, the new input prompts can be constructed by new candidate triggers $\delta_{i}$ along with input queries $x_{i}$, fromally expressed as $p_{ij}=\{I;x_{i}+\delta_{j}\}$, where $\delta_{j}\in B$ and $i\in[0,N]$. Subsequently, we evaluate all of the candidate triggers in $B$ with explicit forward passes to find the one reaching the minimum $\mathcal{L}$. This allows an efficient approximation of the true greedy selection. Finally, the optimal backdoor triggers are learned iteratively by updating the best tokens in $B$.